The equation is "one half of a number increased by 16 is four less than two thirds of the number. x/2 + 16 = 2/3 - 4 We know the answer is x= 120 and the solution is 76 but just don't know how to show the work????
Answer by rfer(16322)

1/2x+16=2/3x-4
3/6x-4/6x=-20
-1/6x=-20
x=-20/1*-6/1
x=120/1=120